1. Consider an economy with a single firm and a single consumer. The firm marginal abatement costs of pollution x, MAC = 120 − .4x, and the consumer suffers marginal damages of pollution MD = .25x.

a. If the firm has the right to pollute, what is the minimum amount of money it would accept from the consumer to abate 50 units of emissions?

b. If the firm has the right to pollute, what is the maximum amount of money the consumer would pay the firm for it to abate 50 units of emissions?

c. If the consumer has the right to a clean environment, what is the minimum amount they would be willing to accept to allow the firm to produce 100 units of emissions?

d. If the consumer has the right to a clean environment, explain the transaction (including payment amounts) that would get the economy to the efficient level of pollution?
Answer #1

a)

In absence of any regulation, firm will emit such that

MAC=0

120-0.4x=0

x=300 units

Firm will emit 300 units in absence of any regulation.

Firm would accept a minimum amout which is equal to total cost of abating 50 units of emissions.

Total Cost of abatement, TAC is given as

Firm will accept a minimum of $500 for that purpose.

b)

Maximum amount payable will be equal to reduction in damages i,e,

c)

Minimum amount acceptable by consumers is damage incurred by firm.

d)

Efficient allocation requires that

MAC=MD

120-0.4x=0.25x

x=184.62

MD=0.25x=0.25*184.62=$46.15

MAC=120-0.4x=120-0.4*184.62=$46.15

So, A tax equal to MD i.e. $46.15 per unit of emission will bring the efficient level of pollution.